Isoscles triangle ABC has vertices A(-1,1), B(1,5), and C(3,1) What is the equation of the perpendicular bisector of line segment BC?
Answer by stanbon(75887)

Isoscles triangle ABC has vertices A(-1,1), B(1,5), and C(3,1) What is the equation of the perpendicular bisector of line segment BC?
------------------
The midpoint of BC is x = (1+3)/2 ; y = (5+1)/2 or (2,3)
-----
BC has slope (5-1)/(1-3) = -2
----
So the perpendicular bisector is the line
with slope 1/2 passing thru point (2,3)
-------
Form y = mx + b
Substitute for y,x, and m, and solve for "b":
-------
3 = (1/2)2 + b
3 = 1 + b
b = 2
------------
Equation:
y = (1/2)x + 2
========================
Cheers,
Stan H.
=============